# Dream to Manipulate: Compositional World Models Empowering Robot Imitation Learning with Imagination

### v1

@@ 3.3 Modelling Robot-Object Dynamics with Object Assets
To be able to predict the consequences of the actions by the robot, we must model the dynamics $D$. We can directly use external physics engines (Todorov et al., 2012; Coumans and Bai, 2016) for the dynamics operator $D$. For our model, we rely on PyBullet (Coumans and Bai, 2016) for its simplicity.
While object-centric Gaussian Splat representations $\mathcal{G}_{k}$ are good for photorealistic renderings of objects, they are not designed for physics simulations. Instead, physics engines —including PyBullet— typically operate on object meshes. We thus need to convert $\mathcal{G}_{k}$ into object mesh grids, $M_{k}=\big{(}\mu_{k,\cdot},\rho_{k,\cdot},u_{k,\cdot}\ \big{)}$, with mesh centers $\mu_{k,\cdot}$, orientations $\rho_{k,\cdot}$, and appearances $u_{k,\cdot}$, to obtain our world object assets, $o_{k}=\big{(}\mathcal{G}_{k},M_{k},\omega\big{)}$. With $\omega$ we explicitly denote the relevant physical parameters of the world, such as the masses of objects or the friction coefficient. In the robot simulator, these are in the URDF files (Villasevil et al., 2024). During learning the policies, the parameters $\omega$ can be either given or inferred (Memmel et al., 2024).
The physics engine acts as an operator on the positional and orientation parameters of the mesh grid of all objects in the world, that is
$\begin{pmatrix}\mu_{\cdot,t+1}\\ \rho_{\cdot,t+1}\end{pmatrix}=\begin{pmatrix}\mu_{\cdot,t}+\Delta\mu_{1\dots K}\\ \Delta\rho_{1\dots K}\cdot\rho_{\cdot,t}\end{pmatrix}$ (1)
@@
So far we assumed that the physical parameters are fixed. With this assumption, we cannot guarantee that the manipulation world model will lead to correct predictions for arbitrary actions, as errors in the Gaussian Splat reconstructions and their mesh grids, or misalignment in the physical parameters, could result in cumulative errors when applying the dynamics. That said, our undertaking in this work is that by adopting explicitly grounded appearance and dynamics models for the world and its objects, we can still obtain a useful model. Suppose that the model predictions for replaying the sequence of actions in the demonstration result in a final state similar to the final state of the demonstration. In that case, we can still use the world model as an interactive and compositional representation of the demonstration itself. In the next section, we show that such interactive representation allows us to augment the original demonstrations with imagined ones to train a more robust imitation learning policy.
@@ 3.4 Engineering
The proposed compositional manipulation world model is a significant undertaking, requiring extensive engineering to ultimately be effective with real-world robots. However, as we demonstrate later, one of the key advantages of this approach is that it enables a real robot to learn novel policies in a one-shot manner, using only a single example. For transparency and reproducibility, we describe the most critical engineering components in detail, with additional information provided in App. J.
While Gaussian Splatting is primarily designed for novel view synthesis, we observed unexpected behavior at low resolutions, leading to incorrect action predictions by the robot. Additionally, we found that the depth information rendered along the edges of objects and the robot was inaccurate, as the original Gaussian Splatting method does not ensure that Gaussians align precisely with surfaces. To address this, we implemented 2DGS Gaussian Splatting (Huang et al., 2024). For successful real-world deployment of DreMa, it is crucial that the geometry of the reconstruction and the simulator are properly aligned, which we achieved using a calibrated camera mounted on the robot (Allegro et al., 2024). Also, estimating physical parameters is a complex field in its own right (Mehra, 1974; Chebotar et al., 2019; Gao et al., 2022; Huang et al., 2023), and methods such as ASID (Memmel et al., 2024) or AdaptSim (Ren et al., 2023) can be applied for this purpose. In this paper, for simplicity we pick constant physical parameters that seemed to work well in the validation sets.
@@ 4 Equivariant Transformations for Imitation Learning
Finding equivariant transformations for high-dimensional observations $\mathbf{Q}$ and actions $\mathbf{A}$ is typically challenging, but DreMa simplifies this by using explicit representations for object assets $\mathcal{O}$. In our world model, both the robot’s actions $\mathbf{A}$ (the 3D positions of the end-effector and gripper) and the 3D positions of the objects, including initial and target locations, are transformed equivariantly under any rigid transformation. As a result, these transformations preserve task success in the imagined demonstrations. This allows us to generate novel “imaginations” by applying geometric transformations to the objects’ positions and the robot’s actions, without changing the robot’s internal state. By executing these transformed actions $\hat{\mathbf{A}}$ and rendering the corresponding observations $\hat{\mathbf{Q}}$ in DreMa, we generate new demonstrations $\hat{\zeta}\in\mathcal{Z}_{new}$ not present in the original dataset $\mathcal{Z}$. Below, we describe several such transformations applicable to manipulation tasks (see Fig. 3).
@@
This transformation involves rotating the objects around the final position of the action sequence $\mathbf{A}$. The transformation $R$ is applied to both the object poses and the action trajectory $\mathbf{A}$, keeping the demonstration consistent with the original task. In this second transformation we fixed $P_{\kappa}=et_{t}$ (the last position of the end-effector). We then rotate the objects assets $o_{k}$ at time $t=0$ and the actions $a_{i}$ similarly to the previous approach. In this transformation we don’t change the environment since the trajectory brings the object to a similar position.
@@ 5.1 One-shot Policy Learning: Single- and Multi-task
We replicate the experimental setup of PerAct (Shridhar et al., 2023) using RLBench (James et al., 2020), which involves a Franka Emika Panda robot with a parallel jaw gripper, placed on a table with objects that vary in color and position. We focus on tasks excluding articulated objects, as our current object-centric priors do not handle articulations. The experiments cover 1) non-prehensile tasks and 2) pick-and-place tasks. Following Ze et al. (2023), we collect a sequence, $\mathbf{X}$, of $200$ images by simulating a rotating camera around the scene. PerAct supports multi-task training, where a single policy adapts to multiple tasks based on prompts. We train both PerAct and DreMa on three tasks using the same demonstrations as in the single-task setup. Hyperparameters are consistent with the original paper (Shridhar et al., 2023), except for the training iterations and batch size (see App. L). To prevent overfitting, we validate all models performance on a separate set for each 10k training iterations for multi-task and 5k for single-task and select the best-performing model.
Table 1: Comparison of PerAct (Shridhar et al., 2023) trained on original demonstrations, invariant augmentations, and DreMa with only imagined demonstrations and also the original. The table reports the mean $\pm$ std and maximum success rate over 5 test runs on 50 test environments.
@@
Since the limited number of demonstrations is insufficient to learn a general policy, the PerAct model trained solely on original demonstrations performs the worst. In contrast, DreMa significantly, exploiting around 800 demonstrations (see App. L), outperforms PerAct with the original data in single-task settings (Table 1), achieving an average accuracy improvement of $+9.1\%$. This result highlights that additional imagination-based demonstrations help create more robust policies, even when the original dataset is small. When DreMa is compared to invariant augmentations —namely, random patches, random table color, and random distractors— it achieves the best performance, with a 5.0% average improvement over the best invariant augmentation (see App. A). Additionally, DreMa improves performance across all tasks compared to PerAct, differently to the invariant augmentations. In multi-task settings (Table A.5), DreMa is tested on close jar, slide block, and sort shape simultaneously. It achieves an average accuracy of $35.5\%$, representing a $+13.1\%$ improvement over PerAct, demonstrating that even this setting can significantly benefit from imagination-based demonstrations. DreMa trained on only imagination demonstrations and evaluated in the original environment still performs significantly better than PerAct ($+3.1$ on single task $+9.1$ on multi-task).
@@ 5.2 Analysis and ablations
We analyze the effectiveness of DreMa ’s imagined demonstrations based on the number of original demonstrations provided. Specifically, we vary the dataset size $|\mathcal{D}|\in\{4,10,20\}$ for the slide block and and $|\mathcal{D}|\in\{5,10,20\}$ sort shape tasks, as they performed the best and worst, respectively, in the previous experiment (see App. L for training details). Figure 4 shows the average success rate of DreMa and PerAct as a function of the original dataset size. Increasing the training examples to 20 improves performance in both tasks. PerAct trained on 20 examples matches DreMa trained with imagined data from just 5 demonstrations (81 and 61 imagined demos for sort shape and slide block, respectively). However, when using the same 20 examples to generate imagined data (437 and 391 imagined demonstrations for sort shape and slide block, respectively), the model improves by 6.4% and 10.4%. This shows that generating novel configurations and actions remains beneficial, even with more demonstrations. Figure C.6 in the appendix shows how augmented data covers a wider portion of the robot’s workspace, enabling the agent to handle configurations beyond those in the original demonstrations.
@@
Table 2: Performance of DreMa trained on single-task demonstrations from different types of transformations. The table reports the mean $\pm$ std and maximum success rate over 5 test runs on 50 examples.
@@ 5.3 Experiments with a Real Robot
In distr. OOD In distr. OOD In distr. OOD In distr. OOD In distr. OOD Avg. all Avg. OOD
@@
Figure 5: Original (top) and imagined demonstration (bottom) after a $90^{\circ}$ rotation transformation
@@ 6 Conclusions
Limitations. Although DreMa demonstrates several strengths, it has limitations. First, it requires full observability of the environment to model physics and appearances. Second, it currently handles only simple objects, but recent advances in automatic URDF estimation could address this issue (Chen et al., 2024b; Lewis et al., 2024). Third, we construct it with an open-loop approach without feedbacks; incorporating them into the world model construction could improve both prediction accuracy and policy learning. Finally, scaling the world model to large and diverse environments may increase simulation complexity. Overall, we believe that future research can address these issues, particularly the second and the third.
@@ Appendix B Imagination results in the real experiment
In Table B.6, we present the complete results for the pick block task in the real-world experiment. This experiment aimed to verify whether the robot can learn purely through imagination, as demonstrated in the simulation results. We observe that with only four original examples, the learned policy performed worse than the initial policy. However, its performance significantly improved as more original examples were added.
@@ Appendix C Data Generation Analysis
As mentioned in the work, a common problem of imitation learning, and particularly behaviour cloning, is generalizing outside the given configuration of the environment. The ability of DreMa to generate high-quality data by imagining new configurations of the environment allows us to increase the training distribution. Figure C.6 shows this effect by plotting the location of the keypoints (robot actions) in the original demonstration and in the generated imagination. This results in improved and more robust policies. Gaussian Splatting guarantees that the domain gap is small while the similar allows to recreate of reasonable interaction between the robot and the objects.
@@ Appendix E Explicit World Models and Learnable Digital Twins
The representation function of the world relies on (i) the Gaussian Splat Object Assets, and (ii) the physics simulator, to generate, aka ”imagine”, new RGB-D representations of novel future world states never seen before. We can use the imagined states to train our policies. This aligns with Ha and Schmidhuber’s idea of an agent being trained entirely within "its own hallucinated dream" [Ha and Schmidhuber, 2018].
Similarly, in our experiments, the PerAct agent was trained using imagined sensory data (images rendered via Gaussian Splatting). Results in Table 1 demonstrate the agent’s ability to leverage this generated data, while Table 4 highlights its effectiveness in a real-world setting. By integrating this generated data with original data, we demonstrate improved performance, further validating the predictive power of the constructed world model.
To summarize, our approach satisfies key components of an explicit world model:
@@
We further highlight the connection between world models and learnable digital twins, as in  Abou-Chakra et al. [2024]. When an agent learns a digital twin from observations and exploits it for decision-making or policy learning, the boundary between the two becomes blurred. Learnable digital twins can be seen as explicit models of the environment. While standard world models are usually implicit, they and more explicit digital twins serve a shared purpose of enabling agents to simulate and reason about their surroundings. This connection is further supported by Yang et al. [2023], which uses implicit and interactive world models as simulators.
@@ Appendix F Tasks descriptions
Close jar Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Lift
@@ F.1 Simulation tasks
Pick a target cup from two colored variations and lift it. The task is successful when the correct cup is lifted.
@@ F.2 Real world tasks
The task consists of picking an object selected by the user and put it inside a box. In each scene, there is a tape, a stapler and a yellow duck (that is a distractor). All the shapes are visible in the scene. The task is considered correct if the robot releases the right object inside the red box. Training demonstrations: The objects are randomly placed in the scene, while box is always on the left side of the robot. Test out of domain: the objects are randomly placed in the scene, but the box is positioned on the other side or in a position far from the training one.
@@ Appendix G Extending the world model to articulated objects
Automatically modeling articulated objects remains an open challenge [Kerr et al., 2024, Liu et al., 2023, Weng et al., 2024]. While DreMa is not currently capable of identifying articulated objects, it supports simulating them—for instance, the robotic arm itself is an articulated object. To model articulated objects, it is primarily necessary to: (1) recognize the parts of the object, (2) identify the type and parameters of joints (e.g., rotational about a point or translational along an axis), and (3) locate the joint on the object. Once these elements are determined, a URDF can be generated for simulation [Chen et al., 2024b].
We believe there are promising directions to address these challenges. For example, a robot’s trajectory could help infer object movement and joint characteristics, while temporal supervision of Gaussian Splatting might enhance part and joint estimation, as recently explored by Kerr et al. [2024]. Another approach involves estimating the articulated structure directly from semantics, as discussed by Chen et al. [2024b]. Currently, we manually segment robot links, but this process could be automated in future iterations using a trained model.
In summary, while our current method does not address articulated objects, we acknowledge their importance and see this as a key direction for future work.
Appendix H Discussion on failure sources.
The current implementation of DreMa is highly dependent on the segmentation mask. As shown in Figure H.10, if the segmentation mask for the objects is incorrect, the robot is unable to interact with them. Similar issues may arise from incorrect mesh predictions, which PyBullet uses to detect collisions. Inaccurate meshes can lead to unexpected collisions or missed collisions.
The verification process avoids generating incomplete or incorrect demonstrations. Incomplete demonstrations can result from singularities in the trajectory or from trajectories requiring robot configurations that are unattainable. Incorrect demonstrations often occur when object rotations move them outside the table, causing them to fall.
When PerAct is trained with such faulty demonstrations, its performance is significantly impacted. Even a small number of problematic demonstrations can render the model incapable of executing the task.
Figure H.10: From top to bottom consequences of wrong segmentation, consequences of wrong mesh prediction, consequence of no check of correct execution. In the first two rows the robot cannot pick the objects, while in the last row, it can not perform the actions they requires unreachable configurations.

### v2

@@ 3.3 Modelling Robot-Object Dynamics with Object Assets
To be able to predict the consequences of the actions of the robot, we must model the dynamics $D$. We can directly use external physics engines (Todorov et al., 2012; Coumans and Bai, 2016) for the dynamic operator $D$. For our model, we rely on PyBullet (Coumans and Bai, 2016) for its simplicity.
While object-centric Gaussian Splat representations $\mathcal{G}_{k}$ are good for photorealistic renderings of objects, they are not designed for physics simulations. Instead, physics engines —including PyBullet— typically operate on object meshes. We thus need to convert $\mathcal{G}_{k}$ into object mesh grids, $M_{k}=\big{(}\mu_{k,\cdot},\rho_{k,\cdot},u_{k,\cdot}\ \big{)}$, with mesh centers $\mu_{k,\cdot}$, orientations $\rho_{k,\cdot}$, and appearances $u_{k,\cdot}$, to obtain our world object assets, $o_{k}=\big{(}\mathcal{G}_{k},M_{k},\omega\big{)}$. With $\omega$ we explicitly denote the relevant physical parameters of the world, such as the masses of objects or the friction coefficient. In the robot simulator, these are in the URDF files (Villasevil et al., 2024). During learning the policies, the parameters $\omega$ can be either given or inferred (Memmel et al., 2024). The physics engine acts as an operator on the positional and orientation parameters of the mesh grid of all objects in the world, that is
$\begin{pmatrix}\mu_{\cdot,t+1}\\ \rho_{\cdot,t+1}\end{pmatrix}=\begin{pmatrix}\mu_{\cdot,t}+\Delta\mu_{1\dots K}\\ \Delta\rho_{1\dots K}\cdot\rho_{\cdot,t}\end{pmatrix}\,,$ (1)
@@
So far we assumed that the physical parameters are fixed. With this assumption, we cannot guarantee that the manipulation world model will lead to correct predictions for arbitrary actions, as errors in the Gaussian Splat reconstructions and their mesh grids, or misalignment in the physical parameters, could result in cumulative errors when applying the dynamics. That said, our undertaking in this work is that by adopting explicitly grounded appearance and dynamics models for the world and its objects, we can still obtain a useful model. Suppose that the model predictions for replaying the sequence of actions in the demonstration result in a final state similar to the final state of the demonstration. In that case, we can still use the world model as an interactive and compositional representation of the demonstration itself. In the next section, we show that such an interactive representation allows us to augment the original demonstrations with imagined ones to train a more robust imitation learning policy.
@@ 3.4 Engineering
The proposed compositional manipulation world model is a significant undertaking, requiring extensive engineering to ultimately be effective with real-world robots. However, one of the key advantages of this approach is that it enables a real robot to learn novel policies using only a single example. We describe the most critical engineering components in detail, with additional information provided in App. J. While Gaussian Splatting is primarily designed for novel view synthesis, we observed unexpected behavior at low resolutions, leading to incorrect action predictions by the robot. Additionally, depth information rendered was inaccurate, as the original Gaussian Splatting method does not ensure that Gaussians align with surfaces. To address this, we implemented 2DGS Gaussian Splatting (Huang et al., 2024). For successful real-world deployment of DreMa, properly aligning the reconstruction and the simulator is crucial, which we achieved using a calibrated camera mounted on the robot (Allegro et al., 2024). Also, estimating physical parameters is a complex field in its own right (Mehra, 1974; Chebotar et al., 2019; Gao et al., 2022; Huang et al., 2023), and methods such as ASID (Memmel et al., 2024) or AdaptSim (Ren et al., 2023) can be applied for this purpose. In this paper, for simplicity we pick constant physical parameters that seemed to work well in the validation sets.
@@ 4 Equivariant Transformations for Imitation Learning
Finding equivariant transformations for high-dimensional observations $\mathbf{Q}$ and actions $\mathbf{A}$ is typically challenging, but DreMa simplifies this by using explicit representations for object assets $\mathcal{O}$. In our world model, both the robot’s actions $\mathbf{A}$ (the 3D positions of the end-effector and gripper) and the 3D positions of the objects, including initial and target locations, are transformed equivariantly under any rigid transformation. As a result, these transformations preserve task success in the imagined demonstrations. This allows us to generate novel “imaginations” by applying geometric transformations to the objects’ positions and the robot’s actions, without changing the robot’s internal state. By executing these transformed actions $\hat{\mathbf{A}}$ and rendering the corresponding observations $\hat{\mathbf{Q}}$ in DreMa, we generate new demonstrations $\hat{\zeta}\in\mathcal{Z}_{new}$ not present in the original dataset $\mathcal{Z}$. In the following, we describe several such transformations applicable to manipulation tasks (see Fig. 3).
@@
This transformation involves rotating the objects around the final position of the action sequence $\mathbf{A}$. The transformation $R$ is applied to both the object poses and the action trajectory $\mathbf{A}$, keeping the demonstration consistent with the original task. In this second transformation we fixed $P_{\kappa}=et_{t}$ (the last position of the end-effector). We then rotate the objects assets $o_{k}$ at time $t=0$ and the actions $a_{i}$ similarly to the previous approach. In this transformation we do not change the environment since the trajectory brings the object to a similar position.
@@ 5.1 One-shot Policy Learning: Single- and Multi-task
We replicate the experimental setup of PerAct (Shridhar et al., 2023) using RLBench (James et al., 2020), which involves a Franka Emika Panda robot with a parallel jaw gripper, placed on a table with objects that vary in color and position. We focus on tasks excluding articulated objects, as our current object-centric priors do not handle articulations. The experiments cover 1) non-prehensile tasks and 2) pick-and-place tasks (nine tasks in total). Following Ze et al. (2023), we collect a sequence, $\mathbf{X}$, of $200$ images by simulating a rotating camera around the scene. PerAct supports multi-task training, where a single policy adapts to multiple tasks based on prompts. We train both PerAct and DreMa on three tasks using the same demonstrations as in the single-task setup. Hyperparameters are consistent with the original paper (Shridhar et al., 2023), except for the training iterations and batch size (see App. L). To prevent overfitting, we validate all models performance on a separate set for each 10k training iterations for multi-task and 5k for single-task and select the best-performing model.
Table 1: Comparison of trained on original, augmented, and DreMa with imagined and all demonstrations, reporting mean ± std and max success over 5 runs on 50 test environments.
@@
Since the limited number of demonstrations is insufficient to learn a general policy, the PerAct model trained solely on original demonstrations performs the worst. In contrast, DreMa, exploiting around 800 demonstrations (see App. L), significantly outperforms PerAct with the original data in single-task settings (Table 1), achieving an average accuracy improvement of $+9.1\%$. This result highlights that additional imagination-based demonstrations help create more robust policies, even when the original dataset is small. When DreMa is compared to invariant augmentations —namely, random patches, random table color, and random distractors— it achieves the best performance, with a 5.0% average improvement over the best invariant augmentation (see App. A). Additionally, DreMa improves performance in all tasks compared to PerAct, differently from invariant augmentations. In multi-task settings (Table A.5), DreMa is tested on close jar, slide block, and sort shape simultaneously. It achieves an average accuracy of $35.5\%$, representing a $+13.1\%$ improvement over PerAct, demonstrating that even this setting can significantly benefit from imagination-based demonstrations. DreMa trained on only imagination demonstrations and evaluated in the original environment still performs significantly better than PerAct ($+3.1$ on single task $+9.1$ on multi-task).
@@ 5.2 Analysis and ablations
We analyze the effectiveness of DreMa’s imagined demonstrations based on the number of original demonstrations provided. Specifically, we vary the dataset size $|\mathcal{D}|\in\{4,10,20\}$ for the slide block and and $|\mathcal{D}|\in\{5,10,20\}$ sort shape tasks (see App. L for training details). Figure 4 shows the average success rate of DreMa and PerAct as a function of the original dataset size. Increasing the training examples to 20 improves performance in both tasks. PerAct trained on 20 examples matches DreMa trained with imagined data from just 5 demonstrations (81 and 61 imagined demos for sort shape and slide block, respectively). However, when using the same 20 examples to generate imagined data (437 and 391 imagined demonstrations for sort shape and slide block, respectively), the model improves by 6.4% and 10.4%. This shows that generating novel configurations and actions remains beneficial, even with more demonstrations. Figure C.6 in the appendix shows how augmented data cover a wider portion of the robot’s workspace, enabling the agent to handle configurations beyond those in the original demonstrations.
@@
Table 2: The mean $\pm$ std and maximum success rate of DreMa trained on single-task demonstrations from different types of transformations over 5 test runs on 50 examples.
@@ 5.3 Experiments with a Real Robot
In distr. OOD In distr. OOD In distr. OOD In distr. OOD In distr. OOD avg. OOD
@@
Figure 5: Original (top) and imagined demonstration (bottom) after a $90^{\circ}$ rotation transformation.
@@ 6 Conclusions
Limitations. Although DreMa demonstrates several strengths, it has limitations. First, it requires full observability of the environment to model physics and appearances. Second, it currently handles only simple objects, but recent advances in automatic URDF estimation could address this issue (see App. G). Third, we construct it with an open-loop approach without feedbacks; incorporating them into the world model construction could improve both prediction accuracy and policy learning. Finally, scaling the world model to large and diverse environments may increase simulation complexity. Overall, we believe that future research can address these issues, particularly the second and the third. Another interesting future direction is to study the effect of DreMa with more recent models (such as RVT2 (Goyal et al., 2024) or Diffusion Policy (Chi et al., 2023)) or approaches that directly exploit RGB images to show the benefit of Gaussian Splatting.
@@ Appendix B Imagination results in the real experiment
In Table B.6, we present the complete results for the pick block task in the real-world experiment. This experiment aimed to verify whether the robot can learn purely through imagination, as demonstrated by the simulation results. We observe that with only four original examples, the learned policy performed worse than the initial policy. However, its performance improved significantly as more original examples were added.
@@ Appendix C Data Generation Analysis
As mentioned in the work, a common problem of imitation learning, and particularly behavior cloning, is generalizing outside the given configuration of the environment. The ability of DreMa to generate high-quality data by imagining new configurations of the environment allows us to increase the training distribution. Figure C.6 shows this effect by plotting the location of the keypoints (robot actions) in the original demonstration and in the generated imagination. This results in better and more robust policies. Gaussian Splatting guarantees that the domain gap is small while the simulator allows us to recreate a reasonable interaction between the robot and the objects.
@@ Appendix E Explicit World Models and Learnable Digital Twins
The representation function of the world relies on (i) the Gaussian Splat Object Assets, and (ii) the physics simulator to generate, aka “imagine”, new RGB-D representations of novel future world states never seen before. We can use the imagined states to train our policies. This aligns with Ha and Schmidhuber’s idea of an agent being trained entirely within “its own hallucinated dream” [Ha and Schmidhuber, 2018].
Similarly, in our experiments, the PerAct agent was trained using imagined sensory data (RGB and depth images rendered via Gaussian Splatting). The results in Table 1 demonstrate the agent’s ability to use this generated data, while Table 4 highlights its effectiveness in a real-world setting. By integrating this generated data with the original data, we demonstrate improved performance, further validating the predictive power of the constructed world model.
In summary, our approach satisfies key components of an explicit world model:
@@
We further highlight the connection between world models and learnable digital twins, as in  Abou-Chakra et al. [2024]. When an agent learns a digital twin from observations and exploits it for decision-making or policy learning, the boundary between the two becomes blurred. Learnable digital twins can be seen as explicit models of the environment. While standard world models are typically implicit, both they and the more explicit digital twins serve the shared purpose of enabling agents to simulate and reason about their surroundings. This connection is further supported by Yang et al. [2023], which uses implicit and interactive world models as simulators.
@@ Appendix F Tasks descriptions
Close jar
@@
Lift Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ F.1 Simulation tasks
Pick a target cup from two color variations and lift it. The task is successful when the correct cup is lifted.
@@ F.2 Real world tasks
The task consists of picking an object selected by the user and put it inside a box. In each scene, there is a tape, a stapler and a yellow duck (that is a distractor). All the shapes are visible in the scene. The task is considered correct if the robot releases the right object inside the brown box. Training demonstrations: The objects are randomly placed in the scene, while box is always on the left side of the robot. Test out of domain: the objects are randomly placed in the scene, but the box is positioned on the other side or in a position far from the training one.
@@ Appendix G Extending the world model to articulated objects
Automatic modeling of articulated objects remains an open challenge [Kerr et al., 2024, Liu et al., 2023, Weng et al., 2024]. While DreMa does not currently identify articulated objects, it does support their simulation (for instance, the robotic arm itself is an articulated object). To model articulated objects, three key steps are required: (1) recognizing the object’s components, (2) identifying joint types and parameters (e.g., rotational about a point or translational along an axis), and (3) determining the joint’s location. Once these aspects are defined, a URDF can be generated for simulation [Chen et al., 2024b].
Several promising directions could help address these challenges. One possible approach would leverages a robot’s or human motion to infer object movement and joint properties, as demonstrated by Kerr et al. [2024] that used foundation models such as SAM [Kirillov et al., 2023] to identify object parts, while an optimization-based approach improves joint movement. Extracting the meshes of the parts and encoding their movement within a URDF would enable the automatic creation of articulated objects for DreMa. Alternatively, the articulated structure could be inferred directly from semantic information, as demonstrated by Chen et al. [2024b], where a model could be trained to reconstruct articulated objects directly.
In the current version, robot links are segmented manually, while the model is assumed to be given. Automating this process could be achieved by segmenting links using a foundation model [Kirillov et al., 2023] or a trained segmentation model for this task.
Appendix H Discussion on failure sources
The current implementation of DreMa is highly dependent on the segmentation mask. As shown in Figure H.10, if the segmentation mask for the objects is incorrect, the robot cannot interact with them. Similar issues may arise from incorrect mesh predictions, which PyBullet uses to detect collisions. Inaccurate meshes can lead to unexpected collisions or missed collisions.
The verification process prevents generating incomplete or incorrect demonstrations. Incomplete demonstrations can result from singularities in the trajectory or from trajectories that require unattainable robot configurations. Incorrect demonstrations often occur when object rotations move them outside the table, causing them to fall.
When PerAct is trained with such faulty demonstrations, its performance is significantly impacted. Even a small number of problematic demonstrations can make the model incapable of executing the task.
Figure H.10: From top to bottom consequences of wrong segmentation, consequences of wrong mesh prediction, consequence of no check of correct execution. In the first two rows the robot cannot pick the objects, while in the last row, it can’t perform the actions they requires unreachable configurations.